OpenStudy (anonymous):

WILL GIVE MEDAL!!!!! Can someone please help me out with algebra questions, im not the best at algebra, and I need to pass.

OpenStudy (anonymous):

WIll GIVE MEDAL

OpenStudy (anonymous):

What is the value of z in the equation 2z + 7 = −5? −6 −1 1 6 The equation 1.6x − 2.2 = 1.4x can be transformed to form which of the following expressions? 3.0x = 2.2 0.2x = −2.2 3.0x = −2.2 0.2x = 2.2 A system of equations is shown below: n = 3m + 6 n − 2m = 2 What is the solution, in the form (m, n), to the system of equations? (−4, −6) (−5, −9) (2, 6) (1, 8) The system of equations shown below is graphed on a coordinate grid: 3y + x = 4 2y − x = 6 Which statement is true about the coordinates of the point that is the solution to the system of equations? It is (−2, 2) and lies on both lines. It is (−5, 3) and lies on both lines. It is (−5, 3) and does not lie on either line. It is (−2, 2) and does not lie on either line. Which statement is true for the equation 3x − 3x − 2 = −2? It has no solution. It has one solution. It has two solutions. It has infinitely many solutions. A system of equations is shown below: Equation A: 2c = d − 8 Equation B: c = 3d + 8 Which of the following steps should be performed to eliminate variable d first? Multiply equation B by 2. Multiply equation A by 2. Multiply equation B by 3. Multiply equation A by −3.

OpenStudy (anonymous):

1 medal for each question helped with :D please im begging

OpenStudy (anonymous):

1.6

OpenStudy (anonymous):

how i got the answer 2z+7=-5 2z=-5-7 2z=-12 2z=-6

OpenStudy (anonymous):

sorry its negative 6

OpenStudy (anonymous):

2.b

OpenStudy (anonymous):

Thank you :D im working on ur question rn

OpenStudy (anonymous):

do you know anymore @CUTEYPIE137

OpenStudy (anonymous):

3. a

OpenStudy (anonymous):

Thank you :D

OpenStudy (anonymous):

5.d

OpenStudy (anonymous):

4 or 5 is d? @CUTEYPIE137

OpenStudy (anonymous):

6. a

OpenStudy (anonymous):

oh 5 is d

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

4.a

OpenStudy (anonymous):

Thank you so much :D